# ScaleLLM: A Resource-Frugal LLM Serving Framework by Optimizing End-to-End Efficiency

### v1

@@ Abstract
Large language models (LLMs) have surged in popularity and are extensively used in commercial applications, where the efficiency of model serving is crucial for the user experience. Most current research focuses on optimizing individual sub-procedures, e.g. local inference and communication, however, there is no comprehensive framework that provides a holistic system view for optimizing LLM serving in an end-to-end manner. In this work, we conduct a detailed analysis to identify major bottlenecks that impact end-to-end latency in LLM serving systems. Our analysis reveals that a comprehensive LLM serving endpoint must address a series of efficiency bottlenecks that extend beyond LLM inference. We then propose ScaleLLM, an optimized system for resource-efficient LLM serving. Our extensive experiments reveal that with 64 concurrent requests,  ScaleLLM  achieves a 4.3$\times$ speed up over vLLM and outperforms state-of-the-arts with 1.5$\times$ higher throughput [^1] https://tensoropera.ai.
@@ 2 Related Work
Many pre-trained open LLMs have been released since last year, where the most commonly used models including Mixtral 8x7B [4] and Llama-3 [15]). Such open-source models motivate the industry to build public LLM-serving endpoints [14, 2] and empower researchers to work on speeding up the inference speed. FlashAttention [1] is proposed to approximate the attention calculation to reduce memory usage with fast computation. By representing the weights and activations with low-precision data types, Model Quantization [9, 10, 19, 20, 18] is also widely adopted to reduce memory and computation costs.
@@ 5.1 Evaluation Metrics
# of Concurrency Requests: The upper bound of the number of ongoing requests at a single moment.
# of Requests: In order to fulfill system during a elapsed time period, this number set to be $20\times c$ where c is the number of concurrency requests.
@@ 5.2 Serving Performance Evaluation
Figure 6: System latency vs # of concurrency requests.
Figure 7: Throughput vs # of concurrency requests.

### v2

@@ Abstract
Large language models (LLMs) have surged in popularity and are extensively used in commercial applications, where the efficiency of model serving is crucial for the user experience. Most current research focuses on optimizing individual sub-procedures, e.g. local inference and communication, however, there is no comprehensive framework that provides a holistic system view for optimizing LLM serving in an end-to-end manner. In this work, we conduct a detailed analysis to identify major bottlenecks that impact end-to-end latency in LLM serving systems. Our analysis reveals that a comprehensive LLM serving endpoint must address a series of efficiency bottlenecks that extend beyond LLM inference. We then propose ScaleLLM, an optimized system for resource-efficient LLM serving. Our extensive experiments reveal that with 64 concurrent requests,  ScaleLLM  achieves a 4.3$\times$ speed up over vLLM and outperforms state-of-the-arts with 1.5$\times$ higher throughput [^1] https://tensoropera.ai/prod/model/mistralai/ScaleLLM-Mixtral-8x7B.
@@ 2 Related Work
Many pre-trained open LLMs have been released since last year, where the most commonly used models including Mixtral 8x7B [4] and Llama-3 [15]). Such open-source models motivate the industry to build public LLM-serving endpoints [14, 2] and empower researchers to work on speeding up the inference speed. FlashAttention [1] is proposed to approximate the attention calculation to reduce memory usage with fast computation. By representing the weights and activations with low-precision data types, Model Quantization [9, 10] is also widely adopted to reduce memory and computation costs.
@@ 5.1 Evaluation Metrics
# of Concurrent Requests: The upper bound of the number of ongoing requests at a single moment.
# of Requests: In order to fulfill system during a elapsed time period, this number set to be $20\times c$ where c is the number of concurrent requests.
@@ 5.2 Serving Performance Evaluation
Figure 6: System latency vs # of concurrent requests.
Figure 7: Throughput vs # of concurrent requests.